Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 384     Entry time: 11 Oct 2023, 06:22
Author: Abhisakh Sarma 
Group: SPB 
Subject: Shift summary 

X-ray delivery

  • 7.8 keV, up to 30 pulses at 94 kHz, 3 mJ

Optical laser delivery

  • No laser

Achievements/Observations

  • Trying to solve issues

Issues

  • AGIPD's asics interlocked several times. Marcin was informed and he tried to solve the issue remotely but did not work.
  • Marcin observed that the AGIPD's chiller temperature was increasing since afternoon and not going down below -16 degree. Probablly the chiller is no longer at proper state. Macrin was onsite and tried to solve the issue without success. Informed to detector group
  • Difficulties with alignment of direct beam path until NKB
ELOG V3.1.4-7c3fd00